OpenStudy (almightyns):

Find Product. (x+y)(4x-5y)

OpenStudy (lord_box):

When multiplying binomials like these together, we use the FOIL method, that is First, Outer, Inner Last. Basically, each term in one () is multiplied by every term in the other (), here is the general equation:\[(a + b)(c + d) = ac +ad+bc+bd\]

OpenStudy (almightyns):

I don't know how to multiply differennt variables

OpenStudy (lord_box):

Here's an example of that: \[(x)(y) = xy, (2x^2)(3y)=6x^2y\]

OpenStudy (almightyns):

so say I pultiply x(5y) it would be 5xy?

OpenStudy (lord_box):

That's right

OpenStudy (almightyns):

Thanks !!

alones (alones):

@AlmightyNs Look at each step (x+y)(4x+−5y) (x)(4x)+(x)(−5y)+(y)(4x)+(y)(−5y) \(4x^2\)−\(5xy \)+4xy−\(5y^2\) and from that way you should get to your answer..

jimthompson5910 (jim_thompson5910):

You can use FOIL or the box method. I prefer the box method because it allows you to go beyond multiplying just binomials